# Detecting Brittle Decisions for Free: Leveraging Margin Consistency in Deep Robust Classifiers

### v1

@@ 6 Conclusion
This work addresses the question of efficiently estimating local robustness in the $\ell_{p}$ sense at a per-instance level in robust deep neural classifiers in deployment scenarios. We introduce margin consistency as a necessary and sufficient condition to use the logit margin of a deep classifier as a reliable proxy estimation of the input margin for detecting non-robust samples. Our investigation of various robustly trained models shows that they have strong margin consistency, which leads to a high performance of the logit margins in detecting vulnerable samples to adversarial attacks and estimating robust accuracy on arbitrarily large test sets using only a small subset. We also find that margin consistency does not always hold, with some models having a weak correlation between the input margin and the logit margin. In such cases, we show that it is possible to learn to map the feature representation to a better-correlated pseudo-margin that simulates the margin consistency and performs better on vulnerability detection. Finally, we present some limitations of this work, mainly the scope of robustness, the attack-based verification and the impact of neural collapse in terminal phases of training. Beyond its highly practical importance, we see this as a motivation to extend the analysis of robust models and the properties of their feature representations in the context of vulnerability detection.
@@ Appendix A Proof of Theorem 1
Formally, for a finite sample $S$ and non-negative values $\epsilon\geq 0$, $\lambda\geq 0$, we define:
@@
Necessity: Let’s assume that the model is not margin-consistent, i.e., there exist two samples $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ such that $d_{out}(\mathbf{x}_{1})\leq d_{out}(\mathbf{x}_{2})$ and $d_{in}(\mathbf{x}_{1})>d_{in}(\mathbf{x}_{2})$. By taking $S=\{\mathbf{x}_{1},\mathbf{x}_{2}\}$ and $\epsilon=d_{in}(\mathbf{x}_{2})$ we have that $A^{S}_{\epsilon}=\{\mathbf{x}_{2}\}$. However for any $\lambda\geq 0$, if $\mathbf{x}_{2}\in B^{S}_{\lambda}$, then $d_{out}(\mathbf{x}_{1})\leq d_{out}(\mathbf{x}_{2})\leq\lambda$ and so $\mathbf{x}_{1}\in B^{S}_{\lambda}$. Therefore $d_{out}$ does not perfectly separates non-robust samples from robust samples.
Sufficiency: Let’s assume that the model is margin-consistent. Let $S$ be a finite sample and consider a threshold $\epsilon$. Let $\mathbf{x}_{0}$ be the element of the finite set $A^{S}_{\epsilon}$ with maximum $d_{in}(\mathbf{x}_{0})$ and $d_{out}(\mathbf{x}_{0})$. Since the model is margin-consistent, then for $\mathbf{x}\in S$:
@@
This means we have $A^{S}_{\epsilon}=B^{S}_{\lambda_{0}},$, which shows that $d_{out}$ perfectly separates non-robust samples from robust samples. ∎
Appendix B Supporting Materials
B.1 Input Margins Estimation Sanity Check
One way to verify the reliability of our estimated input margins is to compare the robust accuracy measured by AutoAttack at $\epsilon=8/255$ and the proportion of correctly classified test samples with estimated input margins greater than $\epsilon$; both quantities should be approximately equal which happens to be the case –See Fig. 8.
@@ B.1 Input Margins Estimation Sanity Check
Figure 8: Robust Accuracy reported in Robustbench vs Robust Accuracy estimated with the input margins estimations. They are both close, indicating a good estimation of the input margins.
B.2 Detection Performance at Different Values of $\epsilon$
We present in Fig. 9 the performance of the detection for various values of the robustness threshold. We can see that the strong margin consistency allows the logit margin to be a good proxy for detection at various thresholds. Note that below $\epsilon=2/255$ and beyond $\epsilon=16/255$, the ratio of vulnerable points to non-vulnerable points becomes too imbalanced, with little to no positive instances beyond $\epsilon=32/255$.
(a) Model AD10 (Addepalli et al., 2021) on CIFAR10
(b) RA11 (Rade & Moosavi-Dezfooli, 2021) on CIFAR100
Figure 9: Variation of AUROC score for different values of the threshold $\epsilon$.
B.3 Sample Efficient Robust Accuracy Estimation
@@
B.4 Pseudo-margin Learning Setup
@@
B.5 Verification of Equidistance Assumption of the Linear Classifiers
Eq. 3 in Sec. 2.1 shows that we can approximate the margin in the feature space by the logit margin if the classifiers $\mathbf{w_{j}}$ are equidistant, i.e. $\|w_{i}-wj\|=C,\,\forall i,j\in\{1,..,C\}$. For each model, we computed the $\frac{C(C-1)}{2}$ possible values of the distances between pairs of classifiers ($45$ for CIFAR10 and $4950$ for CIFAR100). We confirm this hypothesis for our investigated models in Fig. 13 by plotting the boxplot of the distribution of values. For each model, the values vary only in a small range.
(a) CIFAR10
(b) CIFAR100
Figure 13: Equidistance of classifiers: the boxplot reports the distances’ minimum value, lower quartile (Q1), median, upper quartile (Q3), and maximum value. The interquartile range is $Q3-Q1$ is small enough for most of the models.

### v3

@@ 6 Conclusion
This work addresses the question of efficiently estimating local robustness in the $\ell_{p}$ sense at a per-instance level in robust deep neural classifiers in deployment scenarios. We introduce margin consistency as a necessary and sufficient condition to use the logit margin of a deep classifier as a reliable proxy estimation of the input margin for detecting non-robust samples. Our investigation of various robustly trained models shows that they have high margin consistency, which leads to a high performance of the logit margins in detecting vulnerable samples to adversarial attacks. We also find that margin consistency does not always hold, with some models having a weak correlation between the input margin and the logit margin. In such cases, we show that it is possible to learn to map the feature representation to a better-correlated pseudo-margin that simulates the margin consistency and performs better on vulnerability detection. Finally, we present some limitations of this work, mainly the scope of robustness, the attack-based verification, the impact of neural collapse in terminal phases of training, and vulnerability to adaptive attacks. Beyond its highly practical importance, we see this as a motivation to extend the analysis of robust models and the properties of their feature representations in the context of vulnerability detection.
@@ Appendix A Proof of Theorem 1
Formally, for a finite sample $S$ and nonegative values $\epsilon\geq 0$, $\lambda\geq 0$, we define:
@@
Necessity: We proceed by contraposition and assume that the model is not margin-consistent, i.e., there exist two samples $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ such that $d_{out}(\mathbf{x}_{1})\leq d_{out}(\mathbf{x}_{2})$ and $d_{in}(\mathbf{x}_{1})>d_{in}(\mathbf{x}_{2})$. By taking $S=\{\mathbf{x}_{1},\mathbf{x}_{2}\}$ and $\epsilon=d_{in}(\mathbf{x}_{2})$ we have that $A^{S}_{\epsilon}=\{\mathbf{x}_{2}\}$. However for any $\lambda\geq 0$, if $\mathbf{x}_{2}\in B^{S}_{\lambda}$, then $d_{out}(\mathbf{x}_{1})\leq d_{out}(\mathbf{x}_{2})\leq\lambda$ and so $\mathbf{x}_{1}\in B^{S}_{\lambda}$. Therefore $d_{out}$ does not perfectly separates non-robust samples from robust samples.
Sufficiency: Let’s assume the model is margin-consistent. Let $S$ be a finite sample and consider a threshold $\epsilon$. Let $\mathbf{x}_{0}$ be an element of the finite set $A^{S}_{\epsilon}$ such that $d_{in}(\mathbf{x}_{0})=\max\{d_{in}(x):x\in A^{S}_{\epsilon}\}$ and let $\lambda=d_{out}(\mathbf{x}_{0})$. Since the model is margin-consistent, then for every $\mathbf{x}\in S$:
@@
This means we have $A^{S}_{\epsilon}=B^{S}_{\lambda_{0}}$, which shows that $d_{out}$ perfectly separates non-robust samples from robust samples. ∎
Our formulation of perfect separation using finite samples is not fundamentally necessary, however it avoids dealing with the intricacy of the continuum.
Appendix B Detection Performance with Different Values of $\epsilon$
We present in Fig. 7 the performance of the detection for various values of the robustness threshold. We can see that the high margin consistency allows the logit margin to be a good proxy for detection at various thresholds. Note that below $\epsilon=2/255$ and beyond $\epsilon=16/255$, the ratio of vulnerable points to non-vulnerable points becomes too imbalanced, with little to no positive instances beyond $\epsilon=32/255$.
(a) Model AD10 (Addepalli et al., 2021) on CIFAR10
(b) RA11 (Rade & Moosavi-Dezfooli, 2021) on CIFAR100
Figure 7: Variation of AUROC score for different threshold values $\epsilon$.
Appendix C Equidistance Assumption of the Linear Classifiers
In Eq. 3 in Sec. 2.1, we show that we can approximate the exact feature margin by the logit margin when the classifiers $\mathbf{w}_{k}$ are equidistant, i.e. $\|\mathbf{w}_{i}-\mathbf{w}_{j}\|=C$ whenever $i\neq j$. The results in Table 3 show that using the logit margin instead of the exact minimum feature margin has a negligible effect on the results. However, computing the exact feature margin requires computing the minimum over $K-1$ pairs of scaled logit differences. The approximation provided by the logit margin thereby circumvents the computational overhead of the minimum search, which can take a second instead of just microseconds for inference. This difference can add up to hours at scale, offering scalability when dealing with a large number of classes. We additionally provide boxplots for the $K(K-1)/2$ distances between pairs of classifiers for each model ($45$ for CIFAR10 and $4950$ for CIFAR100) in Fig. 8 and 9.
Figure 8: Equidistance of classifiers for CIFAR10 models. The boxplot reports the distances’ minimum value, lower quartile (Q1), median, upper quartile (Q3), and maximum value.
Figure 9: Equidistance of classifiers for CIFAR100 models. The boxplot reports the distances’ minimum value, lower quartile (Q1), median, upper quartile (Q3), and maximum value.
Model $\bm{\tau}$ Lm $\bm{\tau}$ Fm AUROC Lm AUROC Fm AUPR Lm AUPR Fm FPR95 Lm FPR95 Fm Acc Rob. Acc Architecture
CIFAR10 DI0 0.28 0.32 67.49 70.75 70.91 74.28 82.56 80.57 84.36 41.44 WideResNet-28-4
XU80 0.43 0.45 83.30 84.30 80.50 82.25 83.42 82.44 93.69 63.89 WideResNet-28-10
MR0 0.68 0.70 92.95 93.73 94.92 95.54 29.76 27.43 79.69 39.12 ResNet-18
AL0 0.72 0.74 94.67 95.12 95.98 96.28 24.93 22.21 80.38 40.21 ResNet-18
CU80 0.73 0.75 96.87 97.52 94.42 95.47 17.90 14.77 92.16 67.73 WideResNet-28-10
WA80 0.74 0.76 96.82 97.50 94.33 95.43 17.60 14.42 92.44 67.31 WideResNet-28-10
SE10 0.74 0.74 96.03 96.59 94.66 95.52 19.13 16.85 84.59 55.54 ResNet-18
EN0 0.74 0.76 95.16 95.90 95.07 95.85 24.10 21.22 87.03 49.25 ResNet-50
TR0 0.74 0.77 94.63 95.52 96.13 96.74 30.93 25.73 80.72 42.23 ResNet-18
DS0 0.75 0.75 95.80 95.93 95.08 95.23 24.65 23.34 90.06 56.14 XCiT-S12
MD0 0.75 0.75 95.36 95.29 97.00 97.00 23.23 23.75 81.85 36.91 ResNet-18
ZH0 0.75 0.77 95.86 96.35 95.65 96.14 24.91 21.74 84.92 53.08 WideResNet-34-10
AD10 0.82 0.84 97.86 98.26 97.68 98.09 13.26 11.50 80.24 51.06 ResNet-18
CIFAR100 HE1 0.74 0.74 94.43 94.41 97.39 97.39 30.40 30.40 59.23 28.42 WideResNet-28-10
WU1 0.78 0.79 95.81 95.83 98.00 98.01 23.34 22.55 60.38 28.86 WideResNet-34-10
RE812 0.80 0.80 96.87 96.91 98.30 98.32 18.06 17.49 62.41 32.06 WideResNet-28-10
DS1 0.81 0.81 96.78 96.78 98.30 98.30 19.18 18.46 67.34 32.19 XCiT-S12
CU41 0.82 0.82 97.07 97.09 98.48 98.49 17.21 17.35 64.08 31.65 WideResNet-34-10
CU81 0.83 0.83 97.41 97.43 98.24 98.25 15.62 15.67 73.85 39.18 WideResNet-28-10
RI1 0.83 0.83 96.61 96.62 99.05 99.06 18.14 17.70 53.83 18.95 PreActResNet-18
PA21 0.83 0.83 97.66 97.70 98.82 98.84 13.83 13.80 63.66 31.08 WideResNet-28-10
WA81 0.83 0.83 97.51 97.50 98.28 98.27 14.96 14.86 72.58 38.83 WideResNet-28-10
AD21 0.84 0.84 97.46 97.48 98.92 98.93 16.00 15.36 65.45 27.67 ResNet-18
AD1 0.84 0.84 97.65 97.67 98.99 98.99 13.88 13.34 62.02 27.14 PreActResNet-18
RE11 0.85 0.84 97.97 97.88 99.05 99.01 13.21 13.36 56.87 28.50 PreActResNet-18
RA11 0.85 0.85 98.01 98.01 99.08 99.08 12.36 12.20 61.50 28.88 PreActResNet-18
Table 3: Correlations between the input margin and the logit margin (Lm) or the exact Feature margin (Fm) and detection scores on CIFAR10 ($\ell_{\infty}$, $\epsilon=8/255$).
Appendix D Logit Margins of Adversarial Examples
Adversarial examples are perturbed samples that are close to the decision boundary. Therefore, we would expect these samples to have a very small logit margin for strongly margin-consistent models. Here, we study the adversarial examples we used to estimate the input margin. In Table 4, we present the $99^{\text{th}}$ percentile of logit margins of adversarial examples and the detection threshold selected to obtain $95\%$ True Positive Rate. We can observe that the values of the adversarial logit margins are significantly smaller than the detection threshold, so they would be detected as non-robust – just like clean samples that lie close to the decision boundary.
Model ID
Logit margin of
Adversarial Examples
($99^{\text{th}}$ percentile)
Threshold of logit margin
at $95\%$ TPR for detection
CIFAR10 AD1 0.01 1.42
DS0 0.61 2.19
DM0 1.64 3.01
MR0 0.04 8.35
MD0 0.06 11.04
SE10 0.06 2.83
PA20 0.14 1.69
HE0 0.07 2.82
TR0 0.07 4.11
CL0 0.03 5.76
EN0 0.13 5.06
AL0 0.03 5.11
AD20 0.03 1.33
CU80 1.32 2.02
RE802 0.13 1.87
ZH0 0.17 2.87
WA80 0.70 2.14
AD10 0.01 1.20
CIFAR100 AD1 0.01 1.42
AD21 0.02 1.44
CU81 0.13 2.09
CU41 0.04 1.95
DS1 0.13 1.36
HE1 0.06 1.81
PA21 0.08 1.56
RA11 0.08 1.77
RE812 0.17 1.73
RE11 0.07 1.40
RI1 0.03 3.42
WA81 0.09 1.59
WU1 0.06 2.02
Table 4: Logits of the adversarial examples are very small compared to the logit margin for the logit margin threshold for the detection at epsilon=8/255. See Table 1 for the specific references on the model ID.
Appendix E Results on ImageNet and $\ell_{2}$-robust models on CIFAR10
Table 5 shows the results with $\ell_{\infty}$-robust models on ImageNet (Deng et al., 2009), a larger dataset, and results on $\ell_{2}$-robust models (only available on CIFAR10 in Robustbench). The results extend well in both situations.
Model ID Kendall tau AUROC AUPR FPR95 Acc Rob. Acc Architecture
CIFAR10 ($\ell_{2}$) DI02 0.66 95.91 93.05 21.94 88.02 66.09 WideResNet-28-4
AU02 0.69 95.59 87.41 19.30 91.08 72.91 ResNet-50
SE102 0.72 97.07 90.99 13.25 89.76 74.41 ResNet-18
WA802 0.74 98.55 93.59 8.00 95.16 83.68 WideResNet-28-10
RI02 0.74 96.82 92.86 13.98 88.67 67.68 PreActResNet-18
EN02 0.75 96.92 92.78 13.25 90.83 69.24 ResNet-50
RO02 0.76 98.19 96.51 10.15 89.05 66.44 WideResNet-28-10
RE802 0.77 98.72 95.02 6.66 91.79 78.80 WideResNet-28-10
RE102 0.79 98.58 95.32 7.04 90.33 75.86 PreActResNet-18
RA102 0.80 99.00 96.78 5.53 90.57 76.15 PreActResNet-18
ImageNet ($\ell_{\infty}$) SSI0 0.63 90.95 90.06 36.47 72.56 48.08 ViT-S + ConvStem
ENI0 0.73 94.95 97.34 25.07 62.56 29.22 ResNet-50
SAI1 0.74 95.00 96.50 24.88 64.02 34.96 ResNet-50
SAI0 0.77 95.16 97.99 29.31 52.92 25.32 ResNet-18
WOI0 0.78 96.28 98.34 25.65 55.62 26.24 ResNet-50
Table 5: Correlations between the input margin and the logit margin and detection scores on CIFAR10 ($\ell_{2}$, $\epsilon=0.5$) and ImageNet ($\ell_{\infty}$, $\epsilon=4/255$, 1000 samples). See Table 1 for the specific references on the model ID.
Appendix F Additional Applications
F.1 Sample Efficient Robust Accuracy Estimation
Margin consistency enables empirical robustness evaluation over an arbitrarily large test set by only attacking a small subset of test samples. For a robustness evaluation at threshold $\epsilon$ (e.g., $\epsilon=8/255$ in $\ell_{\infty}$ norm on CIFAR10 and CIFAR100), we randomly sample a small subset of the test set and determine the threshold $\lambda$ for the logit margin that best corresponds to $\epsilon$. We propose to use this threshold $\lambda$ to estimate the standard robust accuracy by evaluating the proportion of test samples which are correctly classified and whose logit margin is above $\lambda$ (see Algorithm 1).
Algorithm 1 Sample Efficient Robustness Estimation with Margin Consistency
2:  Output: Robust Accuracy Estimation $\mathcal{A}_{r}$
3:  - Select uniformly at random a subset $X_{n}$ of $n$ samples from $X$.
A naive way to set the threshold $\lambda$ at line 6 of Algorithm 1 would be to set it to the detection threshold at $\alpha=95\%$ TPR or $\alpha=90\%$ TPR, but the logit margin threshold could vary from one model to another; therefore a better way is to select it by tuning over values $\alpha\geq 0.80$ that gives the best approximation of the robust accuracy in terms of the absolute error on the small subset $X_{s}$. The same idea allows estimating a model’s vulnerability over a large dataset without the labels.
We show that this leads to an accurate estimation of the robust accuracy of the investigated models evaluated with over $10,000$ by attacking only a random subset of size $500$. Fig. 10 shows the absolute error of the estimation obtained using $500$ samples. As expected, the estimation over the two weakly margin consistent models is not accurate while having a relatively small absolute difference on the strongly margin consistent models.
@@ F.1 Sample Efficient Robust Accuracy Estimation
F.2 Robustness Bias Analysis using the Logit Margin
Robust models often display robustness bias, namely a disparity of robustness across classes. Interestingly, in a strongly margin-consistent model (see Fig. 13, top row), we show that these discrepancies across classes with respect to input margin are reflected in the logit margin. Additionally, the margin consistency remains strong for each class. However, for a weakly margin-consistent model (Fig. 13, bottom row), significant disparities exist between the correlations across classes, making using logit margin as a proxy for input margin problematic.
Figure 13: Robustness bias analysis with the input margin (first column) vs the logit margin (second column) and correlation scores per class (third column). We have a margin-consistent model on the top row, and on the bottom row, we have a weakly margin-consistent model. On the first column (resp. second column), each boxplot represents the distribution of the input margin (resp. logit margin) for the corresponding CIFAR10 class. The dotted blue line indicates the threshold on the first two columns and is the correlation score computed over all classes. The threshold $\lambda$ in the second column is the logit margin threshold at 95% TPR for detection at $\epsilon=8/255$. The robustness bias of the strongly margin consistent model can be detected using the logit margin, unlike the weakly margin consistent model.
Appendix G Pseudo-margin Learning Setup